# DeepTrace: Learning to Optimize Contact Tracing in Epidemic Networks with Graph Neural Networks

### v3

@@ III-B Contact Tracing by Depth-First Search
Now, let us consider enlarging $\mathbb{G}_{N}$ with DFS tracing strategy [34], which is illustrated in Fig. 3. We can categorize all consecutive pairs $(v_{i}^{*},v_{i+1}^{*})$ in $S$ into the following three types:
@@
Theorem 2 implies that when using the DFS tracing strategy to build the contact tracing network, each node in the epidemic network $\mathbb{G}_{N}$ should not be identified as the superspreader more than $\log_{2}(N)$ times as the contact tracing network expands iteratively.
Next, let us compare the performance of both the DFS and BFS contact tracing strategies using two quantifiable metrics (that can be readily measured once the ground truth is given). The first metric is the average error, which is defined by
@@
The average error metric assesses the performance of any digital contact tracing algorithm, accounting for the selected forward contact tracing strategy and potential suboptimality in maximizing likelihood estimates. On the other hand, the first detection time measures how quickly the contact tracing algorithm can zoom in on identifying the most-likely superspreader without a global view of the overall epidemic networks. It is desired that the most-likely superspreader is the actual ground truth even as the epidemic networks size $N$ increases independently of the digital contact tracing effort.
@@ III-C Interpretation as Online Maze Solving
Interestingly, the above technique to decompose (1) into the respective forward and backward contact tracing subproblems can be interpreted as solving a maze where the epidemic network $\mathbb{G}_{N}$ induces an abstract maze topology. Each step taken by the maze solver corresponds to the discovery of a leaf node $v^{n}$ of the rooted tree $\mathbb{G}_{N}$ at the $n$th iteration of the forward contact tracing stage and a corresponding $v_{n}^{*}$ of the backward contact tracing. We can treat two aforementioned sequences $S=(v_{1}^{*},v_{2}^{*},\ldots,v_{N}^{*})$ and $X=(v^{1},v^{2},\ldots,v^{N})$ as two trajectories on two mazes respectively. The trajectory $S$ is to find the desired exit $v^{*}_{N}$ and the trajectory $X$ traverses the graph. However, these two mazes are explored and solved at the same time, i.e., once the maze solver takes a step from $v^{i}$ to $v^{i+1}$, there is a corresponding movement on $S$ from $v_{i}^{*}$ to $v_{i+1}^{*}$. If we assume that there is only one contact tracer, and each time a new node is discovered in the $(i+1)$th iteration, the contact tracer has to travel from $v^{i}$ to $v^{i+1}$ with the distance $\textup{dist}(v^{i},v^{i+1})$. Then, the goal of the contact tracing is to minimize both the total traveling distance $\sum\limits_{i=1}^{n}\textup{dist}(v^{i},v^{i+1})$ of the contact tracer and the first detection time of $v_{N}^{*}$ (i.e., the optimal maximum-likelihood estimate had $\mathbb{G}_{N}$ been given in advance) during the discovery of all infected nodes in $\mathbb{G}_{N}$. The maze solver (i.e., contact tracer) knows neither the complete maze topology induced by $\mathbb{G}_{N}$ nor the target $v_{N}^{*}$. Upon visiting a vertex (i.e., the graph centroid of $\mathbb{G}_{N}$) in this abstract maze for the first time, the maze solver has information about some of its incident edges only. The contact tracer has a map of all the infected nodes visited and edges connecting them, but cannot tell where each leads until $\mathbb{G}_{N}$ is fully traversed.
Forward contact tracing and backward contact tracing are thus analogous to the process of online maze exploration and maze traversal. Let us connect the aforementioned results on forward contact tracing with this abstract maze interpretation: suppose a given abstract maze is a tree, and there is no cost for each newly discovered node, then Theorem 1 illustrates that the BFS traversal yields the shortest path (i.e., shortest sequence of centroids) to solving the maze, and Theorem 2 describes that the DFS traversal yields a sequence of steps with less variance between steps (i.e., consecutive centroids having least hop distances). In fact, if $\mathbb{G}_{N}$ is a tree, this particular abstract maze constructed using graph centroids will be topologically identical to one constructed by using either the rumor centrality or the distance centrality by virtue of the fact that the graph centroid is equivalent to the rumor center and the distance center [17].
As the maze analogy emerges from the network centrality perspective of solving (1), further insights into its theoretical computational complexity can be gained by linking it to the classical topic of online graph exploration, involving online sequential decision-making to determine which destinations can be reached and when [35, 36]. In particular, the corresponding offline problem for online graph exploration is the Traveling Salesman Problem which has NP-hard complexity. This means that finding the optimal solution of (1) in an online setting is, in general, even NP-hard to approximate. The state-of-the-art online algorithm (also based on a generalization of Depth First Search) has a constant competitive ratio of $16$ for planar undirected graphs [35, 36]. This abstract maze-solving perspective can thus open promising avenues for contact tracing, utilizing either low-complexity greedy algorithms or more advanced algorithmic designs.
More importantly, machine learning techniques can offer new insights to learn to optimize this class of problems in a scalable manner. There are recent works in [37, 38] that used GNN to address online graph exploration as a reinforcement learning problem. In solving (1) online with access to local information, the maze solver builds a complete map of the topology of the abstract maze of graph centroids and, in addition to BFS and DFS traversal, can leverage other maze-solving algorithms in the literature, e.g., Tremaux’s algorithms and the A search algorithm [34]. It is possible to adapt different horizons of forward contact tracing steps to adapt the topological space of this abstract maze and to give weights to its edges. There are also new possibilities to construct the underlying abstract maze by considering different measures of network centrality that can inspire novel jointly optimal forward and backward contact tracing strategies to solving (1).
@@ IV DeepTrace: Learn to Optimize Tracing
In general, solving the NP-hard problem in (2) is computationally challenging, especially when the underlying $\mathbb{G}_{N}$ is large and with $N$ increasing throughout the contact tracing process. To address this challenge, we leverage GNNs to propose a scalable contact tracing algorithm, named DeepTrace (see Section IV-B), capitalizing on the above forward and backward contact tracing decomposition. DeepTrace aims to emulate the online graph exploration strategy to maximize the number of previously unseen unique states in a limited number of steps by employing a graph-structured memory to encode the past trajectory of the contact tracer and hidden structures in the spreading model and network topologies.
IV-A Deep Learning on Graphs
As a deep learning model for processing unstructured data, a GNN extracts graph features and information from the input data that are encoded as graphs. The learning mechanism of the GNN is to iteratively aggregate features and information from neighboring nodes for each node in the input graphs. Aggregating information from neighbors is equivalent to a message-passing process among nodes in a graph. Lastly, we update the values of the learning parameters for regression or classification tasks[39, 40]. To effectively capture both the statistical and graph topology attributes inherent in network-structured input data, the training stage of the GNN in semi-supervised machine learning is crucial. Specifically, the equations (3) and (4) can act as node label generators during the construction of fine-tuning datasets for our proposed Algorithm DeepTrace (see Section IV-C2).
To apply GNNs to solve (2), we need to consider two key computational aspects: the underlying epidemic network $\mathbb{G}_{N}$ is unknown to the contact tracer, and the size of $\mathbb{G}_{N}$ can be potentially massive. The role of the GNN is therefore to facilitate the computation of a (possibly suboptimal) solution to (2) when $\mathbb{G}_{N}$ is large. Due to the adaptability of GNNs, we can first generate a training set using small graphs (e.g., tens or hundreds of nodes) that are subgraphs of $\mathbb{G}_{N}$ and augment the descriptors with the structural features for each individual node of these graphs as the input data. The training stage requires that correct labels (that is, exact values of (3) be associated with each of these small graphs. As the size of the graph grows, the computational cost of (3) becomes expensive. We can instead approximate (3) by sampling method to efficiently compute labels for training data. Finally, we train the GNN model with the input subgraph data iteratively to update the neural network hyperparameters, as shown in Fig. 4. We explain this GNN architecture and its training in the following.
@@ IV-A Deep Learning on Graphs
IV-B Algorithm DeepTrace: Pre-training and Fine-tuning GNNs
In this section, we propose to solve the problem (2) for backward contact tracing in the epidemic network $\mathbb{G}_{N}$ by using GraphSage [40], a popular inductive GNN model, which can be adapted to different network size. To obtain the contact tracing network $\mathbb{G}_{N}$, we construct our GNN model using the LSTM [41] aggregator that aggregates information from the neighbors of a node and the element-wise mean [41] to combine the local information of a node with the aggregated information. We define the $l$th layer of the GNN as follows:
@@ IV-B Algorithm DeepTrace: Pre-training and Fine-tuning GNNs
where $\widehat{\mathbf{w}}^{(l)}$ and $\widecheck{\mathbf{w}}^{(l)}$ are the learning parameters in the LSTM aggregators and combination function, respectively. The initial node feature vector is $\mathbf{h}_{v}^{(0)}=[\mathbf{1},\hat{r}(v),\check{r}(v)]^{T}$ for $v\in\mathbb{G}_{N}$, which we discuss in the Section IV-D. We first prove the effectiveness of backward contact tracing with the GNN model (IV-B) in the following result.
@@
Based on the classic result on the learning capability of feedforward neural networks [42], Theorem 3 states that the value of $\mathbb{P}(G_{n}\mid v)$ can be approximated by GNNs with at most $\textup{diam}(\mathbb{G}_{N})+1$ layers within an arbitrary error $\epsilon$. From an algorithm design perspective, Theorem 3 establishes the feasibility of approximating $\mathbb{P}(G_{n}\mid v)$ using GNNs, providing the theoretical foundation for our subsequent approach to learning $\mathbb{P}(G_{n}\mid v)$ with GNNs.
@@
Since obtaining labeled datasets for training (IV-B) is computationally challenging, we employ a two-phase semi-supervised training approach to tune the learning parameters in (IV-B): a pre-training phase followed by a fine-tuning phase. In the pre-training phase, we train the GNN in Deeptrace using a dataset of contact tracing networks with approximations of $\mathbb{P}(G_{n}\mid v)$ as labels for each node. This dataset can be easily obtained, as described in the Section IV-C. This pre-trained GNN exhibits a great enhancement in performance when compared to a GNN with randomly initialized learning parameters. Then, in the fine-tuning phase, we proceed to adjust the learning parameters in the pre-trained model using an additional dataset of contact tracing networks that contain the exact values of $\mathbb{P}(G_{n}\mid v)$ as labels for each node in the networks. This dataset can include special cases of regularly-sized networks, such as $d$-regular graphs [10, 17], as well as a small number of contact tracing networks that have the exact values of $\mathbb{P}(G_{n}\mid v)$ obtained by calculating probabilities of all the permitted permutations in (2) as labels for each node. Acquiring the latter dataset requires more effort and is considered highly valuable. Therefore, it is not wasted on training the GNN from scratch, but rather for fine-tuning the pre-trained model to further enhance its accuracy. This two-phase training process enables us to effectively utilize both the dataset with approximate values of $\mathbb{P}(G_{n}\mid v)$ and the dataset with exact values of $\mathbb{P}(G_{n}\mid v)$. Thus, we can attain a GNN that exhibits enhanced performance while reducing the time and computational resources required for training. We then apply the fine-tuned models to contact tracing graphs of different scales. This leads to our GNN-based learning algorithm, which we call Algorithm DeepTrace as described below.
As contact tracing progresses, Algorithm 1 ultimately identifies the superspreaders for the entire epidemic network. The computational complexity of contact tracing in Algorithm 1 can be decomposed into three main components: sampling complexity, aggregation complexity, and the complexity of the final regression layer. Suppose there are $N$ nodes in the epidemic network. During the sampling procedure, the GNN samples a fixed-size neighborhood for each node. If each node samples $S$ neighbors and this process is repeated for $L$ layers, then the total number of nodes sampled for each node is $O(S^{L})$. Consequently, the sampling complexity for contact tracing the entire epidemic network is $O(N^{2}S^{L})$. For the aggregation procedure, the GNN aggregates information from its neighbors sampled. Assume that the feature dimension is $m$. Since the LSTM aggregator processes the features of the sampled neighbors, the complexity of the aggregation step is $O(mS^{2})$ per node per layer. Thus, the aggregation complexity for contact tracing the entire epidemic network is $O(N^{2}LmS^{2})$. Regarding the final regression layer in the GNN, a fully connected layer is used to map the node embeddings to the target values, i.e., the maximum likelihood of being superspreaders, of the nodes. Given that the output dimension is $1$, the complexity for each node is $O(m)$. Therefore, the complexity of the final regression layer for tracking the contact throughout the epidemic network is $O(N^{2}m)$. Overall, the computational complexity of the contact tracing in Algorithm 1 is $O(N^{2}LmS^{2})$.

### v4

@@ III-B Contact Tracing by Depth-First Search
Let us consider enlarging $\mathbb{G}_{N}$ with DFS tracing strategy [35], illustrated in Fig. 3. We can categorize all consecutive pairs $(v_{i}^{*},v_{i+1}^{*})$ in $S$ into the following three types:
@@
Theorem 2 implies that when using the DFS tracing strategy to build the contact tracing network, each node in the epidemic network $\mathbb{G}_{N}$ should not be identified as the superspreader more than $\log_{2}(N)$ times as the contact tracing network expands iteratively. Next, let us compare the performance of both the DFS and BFS contact tracing strategies using two quantifiable metrics (that can be readily measured once the ground truth is given). The first metric is the average error, which is defined by
@@
The average error metric assesses the performance of any digital contact tracing algorithm, accounting for the selected forward contact tracing strategy and potential suboptimality in maximizing likelihood estimates. On the other hand, the first detection time measures how quickly the contact tracing algorithm can zoom in on identifying the most likely superspreader without a global view of the overall epidemic networks. It is desired that the most likely superspreader is the actual ground truth even as the epidemic network size $N$ increases independently of the digital contact tracing effort.
@@ III-C Interpretation as Online Maze Solving
We may view splitting (1) into forward/backward tracing as a maze-solving process on the epidemic network $\mathbb{G}_{N}$. Each step taken by the maze solver corresponds to the discovery of a leaf node $v^{n}$ of the rooted tree $\mathbb{G}_{N}$ at the $n$th iteration of the forward contact tracing stage and a corresponding $v_{n}^{*}$ of the backward contact tracing. We can treat two aforementioned sequences $S=(v_{1}^{*},v_{2}^{*},\ldots,v_{N}^{*})$ and $X=(v^{1},v^{2},\ldots,v^{N})$ as two trajectories on two mazes respectively. The trajectory $S$ is to find the desired exit $v^{*}_{N}$, and the trajectory $X$ traverses the graph. However, these two mazes are explored and solved at the same time, i.e., once the maze solver takes a step from $v^{i}$ to $v^{i+1}$, there is a corresponding movement on $S$ from $v_{i}^{*}$ to $v_{i+1}^{*}$. If we assume that there is only one contact tracer, and each time a new node is discovered in the $(i+1)$th iteration, the contact tracer has to travel from $v^{i}$ to $v^{i+1}$ with the distance $\textup{dist}(v^{i},v^{i+1})$. Then, the goal of the contact tracing is to minimize both the total traveling distance $\sum\limits_{i=1}^{n}\textup{dist}(v^{i},v^{i+1})$ of the contact tracer and the first detection time of $v_{N}^{*}$ (i.e., the optimal maximum-likelihood estimate had $\mathbb{G}_{N}$ been given in advance) during the discovery of all infected nodes in $\mathbb{G}_{N}$. The maze solver (i.e., contact tracer) knows neither the complete maze topology induced by $\mathbb{G}_{N}$ nor the target $v_{N}^{*}$.
Upon visiting a vertex (i.e., the graph centroid of $\mathbb{G}_{N}$) in this abstract maze for the first time, the maze solver has information about some of its incident edges only. The contact tracer has a map of all the infected nodes visited and edges connecting them but cannot tell where each leads until $\mathbb{G}_{N}$ is fully traversed. Both forward and backward contact tracing mirror the process of online graph exploration or maze traversal. For a tree with no cost to discover new nodes, Theorem 1 shows that BFS finds the shortest path (i.e., sequence of centroids), while Theorem 2 indicates DFS offers minimal step variance. Moreover, if $\mathbb{G}_{N}$ is a tree, these centroids coincide with the rumor and distance centers [17].
Interpreting forward/backward tracing as an online maze-solving problem links (1) to classic online graph exploration [36, 37], whose offline counterpart (Traveling Salesman Problem) is NP-hard. Thus, finding the optimal online solution of (1) is NP-hard. The best-known online algorithm (a DFS variant) achieves a constant competitive ratio of $16$ on planar graphs [36, 37]. This viewpoint suggests both low-complexity greedy approaches and more advanced algorithms for contact tracing. Additionally, machine learning can learn to optimize such problems at scale [38, 39], e.g., via GNN-based reinforcement learning. When solving (1) online with local data, the maze solver constructs a full map of graph centroids and can integrate maze-solving algorithms like A [35]. Forward tracing horizons and edge weights can be adapted, and alternate centrality measures (e.g., protection [40] or vaccine centrality [17], Bonacich’s centrality [41]) can guide novel forward-backward tracing strategies.
@@ IV DeepTrace: Learn to Optimize Tracing
In general, solving the NP-hard problem in (2) is computationally challenging, especially when the underlying $\mathbb{G}_{N}$ is large and with $N$ increasing throughout the contact tracing process. To address this challenge, we leverage GNNs to propose a scalable contact tracing algorithm named DeepTrace (see Section IV-E), capitalizing on the above forward and backward contact tracing decomposition. DeepTrace aims to emulate the online graph exploration strategy to maximize the number of previously unseen unique states in a limited number of steps by employing a graph-structured memory to encode the past trajectory of the contact tracer and hidden structures in the spreading model and network topologies.
IV-A Overview of GNN-based Framework
As a deep learning model for processing unstructured data, a GNN extracts graph features and information from the input data encoded as graphs. The learning mechanism of the GNN is to iteratively aggregate features and information from neighboring nodes for each node in the input graphs. Aggregating information from neighbors is equivalent to a message-passing process among nodes in a graph. Lastly, we update the values of the learning parameters for regression or classification tasks[42, 43]. To effectively capture both the statistical and graph topology attributes inherent in network-structured input data, the training stage of the GNN in semi-supervised learning is crucial. Specifically, the equations (3) and (4) can act as node label generators during the construction of fine-tuning datasets for the proposed algorithm.
To apply GNNs to solve (2), we need to consider two key computational aspects: the underlying epidemic network $\mathbb{G}_{N}$ is unknown to the contact tracer, and the size of $\mathbb{G}_{N}$ can be potentially massive. The role of the GNN is, therefore, to facilitate the computation of a (possibly suboptimal) solution to (2) when $\mathbb{G}_{N}$ is large. Due to the adaptability of GNNs, we can first generate a training set using small graphs (e.g., tens or hundreds of nodes) that are subgraphs of $\mathbb{G}_{N}$ and augment the descriptors with the structural features for each node of these graphs as the input data. The training stage requires that correct labels (that is, exact values of (3) be associated with each of these small graphs. As the size of the graph grows, the computational cost of (3) becomes expensive. We can instead approximate (3) by sampling method to efficiently compute labels for training data. Finally, we train the GNN model with the input subgraph data iteratively to update the neural network hyperparameters, as shown in Fig. 4. We explain GNN architecture and its training in the following section.
@@ IV-A Overview of GNN-based Framework
IV-B Pre-training and Fine-tuning GNNs
The core idea for solving the backward contact tracing problem described in (2) is to predict the value of $\sum_{\sigma\in\Omega(\mathbb{G}_{n}\mid v)}\mathbb{P}(\sigma|v)$ for each $v\in\mathbb{G}_{n}$ using GNNs. This prediction task can be seen as a node label prediction where the label for each node is a continuous value $\mathbb{P}(\mathbb{G}_{n}\mid v)=\sum_{\sigma\in\Omega(\mathbb{G}_{n}\mid v)}\mathbb{P}(\sigma|v)$. In practice, we use the log-likelihood $\log(\mathbb{P}(G_{n}|v))$ as the node label to avoid arithmetic underflow, as the values of likelihood probabilities can be extremely small as the infected network grows. We employ GraphSage [43], a widely-used inductive GNN model adaptable to varying-sized networks. We use an LSTM-based [43] aggregator to gather information from the neighbors of each node, allowing the model to effectively integrate neighboring node features for learning accurate contact tracing. The $l$th layer of the GNN model is defined as follows:
@@ IV-B Pre-training and Fine-tuning GNNs
where $\widehat{\mathbf{w}}^{(l)}$ and $\widecheck{\mathbf{w}}^{(l)}$ are the learning parameters in the LSTM aggregators and combination function, respectively. The initial node feature vector is $\mathbf{h}_{v}^{(0)}=[\mathbf{1},\hat{r}(v),\check{r}(v)]^{T}$ for $v\in\mathbb{G}_{N}$, which we discuss in the Section IV-D. To train the GNN model in DeepTrace during the two-phase training process, we define the loss functions for the pre-training and fine-tuning phases, respectively, as follows:
$L_{p}(\widehat{\mathbf{w}},\widecheck{\mathbf{w}},v\mid v\in G_{n})$
and
$L_{f}(\widehat{\mathbf{w}},\widecheck{\mathbf{w}},v\mid v\in G_{n})$
where ~$\widetilde{\mathbb{P}}(G_{n}\mid v)$ is the approximation of $\mathbb{P}(G_{n}\mid v)$, $\widehat{\mathbf{w}}=(\widehat{\mathbf{w}}^{(1)},\cdots,\widehat{\mathbf{w}}^{(L)})$ and $\widecheck{\mathbf{w}}=(\widecheck{\mathbf{w}}^{(1)},\cdots,\widecheck{\mathbf{w}}^{(L)})$. The effectiveness of backward contact tracing using GNN (IV-B) is proved as follows.
@@
Based on the classical result on the learning capability of feedforward neural networks [44], Theorem 3 states that the value of $\mathbb{P}(\mathbb{G}_{N}\mid v)$ can be approximated by GNNs with at most $\textup{diam}(\mathbb{G}_{N})+1$ layers within an arbitrary error $\epsilon$. From an algorithm design perspective, Theorem 3 establishes the feasibility of approximating $\mathbb{P}(\mathbb{G}_{N}\mid v)$ using GNNs, providing the theoretical foundation for our subsequent approach to learning $\mathbb{P}(\mathbb{G}_{N}\mid v)$ with GNNs.
@@
Since obtaining labeled datasets for training (IV-B) is computationally challenging, we employ a two-phase semi-supervised training approach to tune the learning parameters in (IV-B): a pre-training phase followed by a fine-tuning phase. In the pre-training phase, we train the GNN in Deeptrace using a dataset of contact tracing networks with approximations of $\mathbb{P}(G_{n}|v)$ as labels for each node. This dataset can be easily obtained, as described in Section IV-C. This pre-trained GNN exhibits a great enhancement in performance when compared to a GNN with randomly initialized learning parameters. Then, in the fine-tuning phase, we proceed to adjust the learning parameters in the pre-trained model using an additional dataset of contact tracing networks that contain the exact values of $\mathbb{P}(G_{n}\mid v)$ as labels for each node in the networks. This dataset can include special cases of regularly-sized networks, such as $d$-regular graphs [10, 17], as well as a small number of contact tracing networks that have the exact values of $\mathbb{P}(G_{n}\mid v)$ obtained by calculating probabilities of all the permitted permutations in (2) as labels for each node. Acquiring the latter dataset requires more effort and is considered highly valuable. Therefore, it is not wasted on training the GNN from scratch but rather on fine-tuning the pre-trained model to further enhance its accuracy. This two-phase training process enables us to effectively utilize both the dataset with approximate values of $\mathbb{P}(G_{n}\mid v)$ and the dataset with exact values of $\mathbb{P}(G_{n}\mid v)$. Thus, we can attain a GNN that exhibits enhanced performance while reducing the time and computational resources required for training.
IV-C Data Construction & Annotation for Two-phases Training
[…]
Next, we apply the SI spreading model [10, 23] on these synthetic networks to simulate the spread of infection to generate the epidemic network $\mathbb{G}_{N}$. Initially, we uniformly select a source node to initiate the spread of the virus on $\mathbb{G}$. The probability of a node in the graph boundary $\mathcal{B}(G_{n})$ being infected in the next time period is uniformly distributed among all outreaching edges from the infected nodes. For example, in Fig. 5, had $v_{3}$ been the source, the probability of $v_{1}$, $v_{2}$, $v_{4}$, and $v_{5}$ being the second infected node is $1/4$, since there are $4$ outreaching edges from $v_{3}$. This process continues until the number of infected nodes reaches our predetermined threshold.
IV-C1 Dataset Construction and Annotation for Pre-training
We generate $500$ underlying networks $\mathbb{G}$ with $3500$ nodes, where each $\mathbb{G}$ is uniformly selected from one of the six types of synthetic networks described above. For each $\mathbb{G}$, we construct $\mathbb{G}_{N}$ according to the SI model with $N$ randomly chosen from $50-1000$. For the labeled training data in the pre-training phase, we can use an approximate estimator of (1) to efficiently obtain ~$\widetilde{\mathbb{P}}(G_{n}\mid v)$ for each node $v$ in the epidemic networks, minimizing time and effort. When the epidemic networks are degree-irregular trees, the ML estimator in (1) aims to calculate the sum of all probabilities of permitted permutations. We can rewrite (1) as follows:
$=\arg\max_{v\in G_{n}}\widetilde{\mathbb{P}}(G_{n}\mid v).$
IV-C2 Dataset Construction and Annotation for Fine-tuning
Since we already have a pre-trained GNN model, we can further refine it using a small amount of high-quality data. Specifically, this involves using data from epidemic networks with exact likelihood probabilities assigned to each node. We generate $250$ underlying networks $\mathbb{G}$ with $3500$ nodes, where each $\mathbb{G}$ is uniformly selected from one of the six types of synthetic networks described above. For each $\mathbb{G}$, we construct $\mathbb{G}_{N}$ according to the SI model with a fixed size of $N=50$. To construct a piece of labeled data with exact likelihood probability for fine-tuning, a natural idea is to calculate the likelihood probability of each node directly by using (3) or (4). We first generate an epidemic network $\mathbb{G}_{N}$ with $N$ nodes, then for each node $v\in\mathbb{G}_{N}$ we generate all permitted permutation $\sigma\in\Omega(\mathbb{G}_{N}\mid v)$ starting from $v$ and compute $\sum_{\sigma\in\Omega(G_{n}\mid v)}\mathbb{P}(\sigma|v)$ using (3). Furthermore, we can construct data samples for which exact likelihood probabilities are easily obtainable using (4), such as the $d$-regular tree. In this case, since the probabilities of all permitted permutations are identical, it is evident from (10) that the likelihood probability for a node in a $d$-regular tree, as given by (2), is proportional to the epidemic centrality in [17], which can be computed using the message-passing algorithm described in [17].
Lastly, we repeat the data generation procedure in the pertaining phase to generate $250$ underlying networks and corresponding epidemic networks for the testing set.
IV-D Node Features Construction
To ensure the effectiveness of GNN learning for solving (2) in Algorithm DeepTrace, it is crucial to design the node features meticulously. This careful design enables the extraction of node information from the underlying epidemic network $\mathbb{G}_{N}$, corresponding to a specific spreading model during the training phase. We now consider several node features, such as the proportion of instantaneously infected nodes and boundary distances within the graph. Notably, constructing these features requires only linear computational complexity, $O(N)$, as illustrated in Fig. 5.
$\mathbf{Infected\;proportion}$: The ratio of the number of infected neighbors of a node $v_{i}$ to the number of all its neighbors:
$\hat{r}(v_{i})=\frac{\hat{d}(v_{i})}{d(v_{i})},$
where $\hat{d}(v_{i})$ is the number of infected neighbors of the node $v_{i}$. For example, the number of infected neighbors of nodes $v_{1},v_{2},v_{3},v_{4},$ and $v_{5}$ in Fig. 5 are $1,1,4,1$ and $1$, leading to an infected proportion of nodes $v_{1},v_{2},v_{3},v_{4},$ and $v_{5}$ being $\frac{1}{2},\frac{1}{3},1,\frac{1}{3},$ and $\frac{1}{4}$, respectively. This feature represents each node’s ability to infect new individuals. Intuitively, a node with a lower infected proportion has a higher probability of infecting new individuals and a smaller probability of being the source. Moreover, this particular node feature aims to capture the effect of the particular node position in the permitted permutation in (3). A node with a lower infected proportion ratio tends to be located at the front of the permitted permutation order, resulting in a smaller permutation probability. For example, as in Fig. 5, the different positions of Node $5$ for the permitted permutation $\{3,1,2,4,5\}$ and $\{5,3,1,2,4\}$ result in a probability of $1.7857\times 10^{-3}$ and $1.0204\times 10^{-3}$, respectively.
$\mathbf{Boundary\;distance\;ratio}$: As the infectivity of the virus evolves–newly symptomatic infectors generally exhibit higher infectivity, whereas long-term symptomatic infectors display lower infectivity–we assume that nodes at the boundary of the epidemic network are newly symptomatic infectors, while those further from the boundary are long-term symptomatic infectors. This characteristic is captured by the boundary distance ratio $\check{r}(\cdot)$. To define $\check{r}(\cdot)$, we first define the graph boundary of $G_{n}$ as $\mathcal{B}(G_{n})=\{u\notin V(G_{n})|\exists v\in G_{n}\text{ s.t. }\textup{dist}(u,v)=1\}$ and the boundary distance $b(v)$ of node $v$ on $G_{n}$ as $b(v)=\min\limits_{u\in\mathcal{B}(G_{n})}\{\textup{dist}(u,v)\}+1$. The boundary distance ratio of $v$ on $G_{n}$ can now be defined as
$\check{r}(v)=\frac{b(v)}{\max_{v_{j}\in G_{n}}b(v_{j})}.$
For example, let $V(G_{n})=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ as shown in Fig. 5, and $\mathcal{B}(G_{n})$ are those unlabeled nodes. Then, we have $b(v_{1})=2$ and $b(v_{3})=3$. Thus, $\check{r}(v_{1})=\frac{2}{3}$, and $\check{r}(v_{3})=1$. Notice that we grow $G_{i}$ using tree traversal algorithms; hence $G_{i}$ can be treated as a tree, and the boundary distance ratio of each node in $G_{i}$ is well-defined. From (2), we see that finding the source also requires information on the size of the collection of all permitted permutations $\Omega(G_{n}\mid v)$. Notice that a node with a larger boundary distance ratio tends to have more permitted permutations. For example, there are $24$ permitted permutations when we choose node $v_{3}$ to be the source, but only $6$ permitted permutations if we choose either node $v_{1},v_{2},v_{4}$ or $v_{5}$ to be the source.
Figure 5: Illustration of the boundary distance for each node on $G_{n}$.
IV-E The DeepTrace Algorithm
Inspired by the general framework discussed in Section IV-A, we divide the DeepTrace algorithm into two main phases: training and prediction. In the training phase, the GNN model is first pre-trained on contact tracing networks where node labels approximate the likelihood function ~$\widetilde{\mathbb{P}}(G_{n}|v)$. This step allows the GNN to gain an initial understanding of networked structures and learn how to approximate the infection source probabilities using these approximated labels. The second step involves fine-tuning the model using the exact values of $\mathbb{P}(G_{n}|v)$. This fine-tuning process refines the predictive capabilities of our model, enabling it to produce more accurate predictions. As the model is trained on precise source probabilities, it adjusts its parameters to optimize performance in the context of source detection. The inputs for training the GNN during the pre-training and fine-tuning phases consist of the corresponding networks that include node features. For testing the trained GNN, the inputs are the contact tracing networks from the testing dataset. The outputs from all phases–pre-training, fine-tuning, and testing are the predicted log-likelihood of each node being the source. We then find the node with the maximum output value as the predicted source.
In the prediction phase, the algorithm begins by generating the contact tracing network $G_{n}$ at the $n$-th stage, using either BFS or DFS. Node features $\mathbf{h}_{v}^{0}$ for each node $v\in G_{n}$ are initialized using $\hat{r}(v)$ and $\check{r}(v)$. These node features are then input into the fine-tuned GNN model, which predicts the likelihood of being the source for each node in $G_{n}$ as shown in Fig. 4. By leveraging the trained GNN model, the algorithm efficiently computes the likelihood of each node being the source of the infection. The prediction phase allows the algorithm to trace the source in the epidemic network.